The area of a town's square is 14,400 square feet. Find the side length of the square.
Found 2 solutions by Alan3354, ReadingBoosters:
Answer by Alan3354(69443)

The area of a town's square is 14,400 square feet. Find the side length of the square.
------------
It's a square.
Side length = s
s*s = 14400
?

Answer by ReadingBoosters(3246)

a = (l)(w)
However, with a square l=w; so,
a = (l)(l) = l^2 = 14400
sqrt(l^2) = sqrt(14400)
l = 120 ft.
.......................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.